OpenStudy (anonymous):

Help please! A glider is released from rest with a spring stretched 0.100 m. What is the displacement, x, of the glider from its equilibrium position when its speed is 0.20 m/s?

OpenStudy (anonymous):

x = 0.1 m v = 0.2 m/s do we know what the spring constant is from the the problem?

OpenStudy (anonymous):

I believe it is 5.00 N/m

OpenStudy (anonymous):

that works for me, so now we know: x = 0.1 m v = 0.2 m/s k = 5 n/m is there any other information, if not, this is a tough one!

OpenStudy (anonymous):

the mass of the glider?

OpenStudy (anonymous):

lets see; m=.200 kg aannndd I think that is it haha

OpenStudy (anonymous):

and that is all we need! is your class studying kinematic equations, momentum or energy?

OpenStudy (anonymous):

umm we just learned conservative energy

OpenStudy (anonymous):

ok great, that is what we will use then -

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

sorry, one last question, is the glider being shoot up into the air? or horizontal?

OpenStudy (anonymous):

horizontal

OpenStudy (anonymous):

so we will break this up in to events event 1 is when the glider is at rest event 2 is when the glider is traveling at a speed of .2 m/s

OpenStudy (anonymous):

ummm okay

OpenStudy (anonymous):

the energy equation will look like this: KE_spring = KE_spring + KE event 1 will have the values of V_1 = 0 m/s K = 5 n/m x = .1 can you remember the kinetic energy equations for KE and KE_spring

OpenStudy (anonymous):

oh and event 1 is on the left of the = sign, and event 2 is on the right of the = sign, btw

OpenStudy (anonymous):

isn't it 1/2mv^2 ?

OpenStudy (anonymous):

or is it 1/2kx^2?

OpenStudy (anonymous):

yep, and for a spring its (1/2) k x^2

OpenStudy (anonymous):

if we exchange this equation KE_spring = KE_spring + KE for this equation (1/2) k x^2 = (1/2) k x^2 + (1/2)mv^2 we just now need to plug in what we know from event 1 and event 2 I'll give you a hint, the x on the left side = 0.1 and we are solving for the x on the right, so we have all the numbers to put in except the x on the right, can you give it a try? i'll draw a picture so it makes a bit more sense while you work on it

OpenStudy (anonymous):

okay thank you! I will see what i get! :)

OpenStudy (anonymous):



OpenStudy (anonymous):

brb

OpenStudy (anonymous):

oh yay i got the right answer! thank you :) so the book says that there are going to be two answers for this problem. one negative and one positive. I don't understand why though?